from the ten books that you've recently bought but not read, you plan to take 4 with you on vacation. how many different sets of 4 books can you take?
Answer by Maths68(1474)

Use
nCr=n!/r!(n-r)!
Here
n=10
r=4
Total number of sets of 4 books = 10!/4!(10-4)!
Total number of sets of 4 books = 10!/4!(6)!
Total number of sets of 4 books = 10*9*8*7*6!/4!(6)!
Cancel out 6!
Total number of sets of 4 books = 10*9*8*7/4!
Total number of sets of 4 books = 10*9*8*7/4*3*2
Total number of sets of 4 books = 5040/24
Total number of sets of 4 books = 210
210 different sets are possible out of 10 books.